OpenStudy (anonymous):

name the property the equation illustrates 7+(4+4)=(7+4)+4...A.)Commutative Property of Addition..B.)Inverse Property of Addition..C.)Associative of Addition..D.)Commutative Property of Multiplicaion

OpenStudy (mathstudent55):

Commutative Property: change the order Associative Property: change the grouping Inverse Property: a + (-a) = 0

OpenStudy (zehanz):

Learn by heart these properties: a+b=b+a (Commutative Property of Addition) (a+b)+c=a+(b+c) (Associative Property of Addition) Then make the right choice.

OpenStudy (anonymous):

it will be Associative property?

OpenStudy (mathstudent55):

Yes

OpenStudy (anonymous):

thank yall so much

OpenStudy (zehanz):

YW!